# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

### v2

@@ 3.3 Enhanced Acquisition Strategy
3.3.1 Weak-to-strong Method
@@
1. Selection of Superior Configurations: From the already evaluated configurations, we select a small number (set to 5 in our experiments) of the top-performing configurations.
2. Differential Evolution: We apply Differential Evolution (DE) [26] to the parameters of these selected configurations. DE is a population-based optimization algorithm that creates new candidate solutions by combining the parameters of existing ones, ensuring diversity and exploration of the search space. Specifically, DE generates new candidate solutions using the following formula:
@@
3. Stochastic Perturbation: To introduce an element of randomness and avoid local optima, we apply a stochastic perturbation with a probability of 50%. This perturbation replaces some of the parameters in the candidate configurations with those obtained from the DE process, encouraging exploration beyond the immediate neighborhood of the superior configurations.
@@
3.3.2 Fisher Information
In the context of our model merging scenario, the purpose is to identify high-quality configurations within a limited number of evaluation iterations. To achieve this, we need to refine our surrogate model efficiently. One way to do this is by carefully selecting which configurations to evaluate based on their potential to improve the surrogate model’s accuracy. Fisher information [27] serves as a pivotal tool in this selection process. By prioritizing configurations with lower Fisher information for evaluation, we can enhance the precision of our surrogate model more efficiently.
@@ 3.3.2 Fisher Information
1. Posterior Predictive Calculation: For each candidate configuration, the Gaussian process surrogate model is employed to compute the posterior predictive mean and variance.
2. Fisher Information Approximation: The variance of the posterior distribution is utilized to approximate the Fisher information for each configuration. Configurations with lower Fisher information indicate regions where the surrogate model’s predictions are less certain.
3. Environment Selection: Configurations with lower Fisher information are prioritized for evaluation. This ensures that configurations contributing more significantly to the model’s uncertainty are evaluated first, thereby enhancing the surrogate model’s overall accuracy.
@@
3.4 Sparsity optimization objective
@@
4 Experimental Setup
4.1 Source Models
To develop a model with strong comprehensive abilities across various disciplines, we applied our proposed model merging method to a set of source models, including Qwen1.5-7B-Chat [29], Liberated-Qwen1.5-7B [30], and firefly-qwen1.5-en-7B [31]. Besides these 7B models, we also experimented with models of different sizes and pre-trained backbones, including Qwen2-1.5B-Instruct [32], SauerkrautLM-1.5b [33], and Qwen2-1.5B-Ita [34]. Please see Table 5 in Section B.2 for their versions and correspondences with pre-trained backbones.
4.2 Datasets
@@
In the configuration search phase, we aimed to optimize the model parameters without overfitting to the test datasets. For this purpose, we used the C-EVAL [22] Validation Set, a Multi-Level Multi-Discipline Chinese Evaluation Suite, to ensure the model’s comprehensive abilities across different disciplines were appropriately calibrated. Additionally, to evaluate and optimize the mathematical reasoning abilities of our model, we sampled 1% of the GSM8K [23] Training Set. Since GSM8K does not have a predefined test set, the validation set is typically used for testing. To avoid data leakage, we selected a small, randomly chosen subset from the training set.
@@
Among these, C-EVAL and MMLU are multidisciplinary examinations that evaluate the model’s comprehensive abilities; GSM8K assesses the model’s mathematical reasoning abilities; HellaSwag evaluates the model’s commonsense reasoning abilities; HumanEval and MBPP assess the model’s code generation abilities; WinoGrande evaluates the model’s language understanding abilities; and BBH assesses the model’s comprehensive reasoning abilities. Please see Section B.3 for more descriptions.
@@
4.3 Baselines
We have compared MM-MO with 5 state-of- the-art and classical merging methods, including Linear (Model Soups) [12], Task Arithmetic [13], TIES [17], DARE [21] and Model Breadcrumbs [41]. Since some model merging methods can be combined, there are a total of 7 baselines. Please see Section B.4 for more descriptions.
4.4 Evaluation Metrics
@@
5 Experimental Results and Analysis
5.1 General Performance
Table 1: Performance Comparison of different merging methods and single models. (Model size: 7B params)
@@ 5.1 General Performance
MM-MO (Ours) Single Model 1 + 2 + 3 60.97 71.9 57.77 74.44 55.49 42.20 60.81 64.17
@@
5.2 Comparison on BBH Dataset
Table 2: Average accuracy over 27 BBH tasks.
@@ 5.2 Comparison on BBH Dataset
MM-MO (Ours) Single Model 1 + 2 + 3 47.72
@@
To validate our methods on difficult tasks, we utilize BBH, which includes a suite of 27 challenging BIG-Bench tasks that require multi-step reasoning. Table 2 shows the performance of single models and the MM-MO merged model on the BBH datasets. Our MM-MO merged model achieves an average BBH score of 47.72, significantly outperforming each of the single models. This result highlights the effectiveness of our model merging strategy in enhancing overall performance.
@@
To further illustrate the advantages of our MM-MO method, we provide a detailed analysis of the performance across various sub-tasks within the BBH dataset. Figure 2 visualizes the scores of single models and the MM-MO merged model on different sub-tasks, As can be seen from the radar chart, MM-MO (red) covers the largest area and has better performance. Notably, in the Temporal Sequences task, MM-MO achieves a score of 47.60, outperforming the best individual model, which scored 40.80. In the Tracking Shuffled Objects (Three Objects) task, MM-MO reaches 54.40, significantly exceeding Single Model 2’s score of 42.00. Furthermore, in the Penguins in a Table task, MM-MO scores 62.33, surpassing Single Model 3’s 56.85. This demonstrates that MM-MO not only leverages the strengths of individual models but also explores their potential further.
5.3 Effectiveness of MM-MO with Different Parameter Scales and Source Models
@@
Table 3: Performance comparison of different merging methods and single models. (Model size: 1.5B params)
@@
Single Model 4 Qwen2-1.5B-Instruct 47.05 68.2 62.09 40.77 42.68 30.00 54.58 42.70 35.37
Single Model 5 SauerkrautLM-1.5b 35.08 38.9 43.44 46.54 23.78 0.00 50.52 51.30 26.18
Single Model 6 Qwen2-1.5B-Ita 46.10 66.9 49.51 48.48 39.63 32.40 53.31 52.41 26.17
MM-MO (Ours) Single Model 4 + 5 + 6 51.04 69.3 63.15 50.65 43.90 35.60 54.84 54.54 36.33
To validate the versatility and generalization capabilities of our method, we applied MM-MO for model merging on source models with different parameter scales and different Pre-Trained Backbones, followed by extensive evaluations of the merged models. As shown in Table 3, the models obtained using MM-MO consistently outperformed any single model across all task types.
Furthermore, it is noteworthy that we selected three models that underwent SFT on different language datasets. Specifically, Qwen2-1.5B-Instruct was fine-tuned on Chinese and English datasets,
SauerkrautLM-1.5b on German, and Qwen2-1.5B-Ita on Italian. MM-MO demonstrated strong adaptability and generalization capabilities in merging models of different languages. This means that even models specializing in different languages can achieve significant performance improvements in multilingual tasks when merged using MM-MO. These results validate the broad applicability of the MM-MO method, showcasing its immense potential for merging models in multilingual, multi-task environments.
5.4 Ablation Study
Table 4: Ablation study results.

### v3

@@ IV-D Enhanced Acquisition Strategy
IV-D1 Weak-to-strong Method
@@
  1. 1.
    Selection of Superior Configurations: From the already evaluated configurations, we select a small number (set to 5 in our experiments) of the top-performing configurations.
  2. 2.
    Differential Evolution: We apply Differential Evolution (DE) [36] to the parameters of these selected configurations. DE is a population-based optimization algorithm that creates new candidate solutions by combining the parameters of existing ones, ensuring diversity and exploration of the search space. Specifically, DE generates new candidate solutions using the following formula:
@@
  3. 3.
    Stochastic Perturbation: To introduce an element of randomness and avoid local optima, we apply a stochastic perturbation with a probability of 50%. This perturbation replaces some of the parameters in the candidate configurations with those obtained from the DE process, encouraging exploration beyond the immediate neighborhood of the superior configurations.
@@
IV-D2 Fisher Information
In the context of our model merging scenario, the purpose is to identify high-quality configurations within limited number of evaluation iterations. To achieve this, we need to refine our surrogate model efficiently. One way to do this is by carefully selecting which configurations to evaluate based on their potential to improve the surrogate model’s accuracy. Fisher information [37] serves as a pivotal tool in this selection process. By prioritizing configurations with lower Fisher information for evaluation, we can enhance the precision of our surrogate model more efficiently.
@@ IV-D2 Fisher Information
  1. 1.
    Posterior Predictive Calculation: For each candidate configuration, the Gaussian process surrogate model is employed to compute the posterior predictive mean and variance.
  2. 2.
    Fisher Information Approximation: The variance of the posterior distribution is utilized to approximate the Fisher information for each configuration. Configurations with lower Fisher information indicate regions where the surrogate model’s predictions are less certain.
  3. 3.
    Environment Selection: Configurations with lower Fisher information are prioritized for evaluation. This ensures that configurations contributing more significantly to the model’s uncertainty are evaluated first, thereby enhancing the surrogate model’s overall accuracy.
@@
IV-E Sparsity Optimization Objective
@@
V Experimental Setup
V-A Source Models
To develop a model with strong comprehensive abilities across various disciplines, we applied our proposed model merging method to a set of source models, including Qwen1.5-7B-Chat [3], Liberated-Qwen1.5-7B [39], and firefly-qwen1.5-en-7B [40]. Besides these Qwen1.5-7B models, we also experimented with models of different sizes and pre-trained backbones (Qwen2-1.5B[41] and Llama-2-13b[6]), including Qwen2-1.5B-Instruct [41], SauerkrautLM-1.5b [42], Qwen2-1.5B-Ita [43], WizardLM-13B [44], WizardMath-13B [45] and llama-2-13b-code-alpaca [46]. Please see Table I for their versions and correspondences with pre-trained backbones.
TABLE I: Versions and correspondences with pre-trained backbones of SFT LLMs.
Parameter Scale SFT LLMs URL Pre-Trained Backbones
13B WizardLM-13B [44] https://huggingface.co/WizardLM/WizardLM-13B-V1.2 Llama-2-13b [6]
WizardMath-13B [45] https://huggingface.co/WizardLM/WizardMath-13B-V1.0
llama-2-13b-code-alpaca [46] https://huggingface.co/layoric/llama-2-13b-code-alpaca
V-B Datasets
@@
In the configuration search phase, we aimed to optimize the configurations of MM without overfitting to the test datasets. For assessing the model’s general abilities, we used the C-EVAL [31] Validation Set, a Multi-Level Multi-Discipline Chinese Evaluation Suite, to ensure the model’s comprehensive abilities across different disciplines were appropriately calibrated. Notably, when merging Llama-series models, we replaced C-EVAL with MMLU [47], considering the Llama models’ stronger proficiency in answering English-based questions. MMLU, which also evaluates general model capabilities, provided a more suitable benchmark for these models. We present the impact of using different language validation sets in Appendix B-E. Additionally, to evaluate and optimize the mathematical reasoning abilities of our model, we sampled 1% of the GSM8K [32] Training Set. Since GSM8K does not have a predefined test set, the validation set is typically used for testing. To avoid data leakage, we selected a small, randomly chosen subset from the training set.
@@
Among these, C-EVAL and MMLU are multidisciplinary examinations that evaluate the model’s comprehensive abilities; GSM8K assesses the model’s mathematical reasoning abilities; HellaSwag evaluates the model’s commonsense reasoning abilities; HumanEval and MBPP assess the model’s code generation abilities; WinoGrande evaluates the model’s language understanding abilities; and BBH assesses the model’s comprehensive reasoning abilities. All detailed descriptions of the datasets are as follows:
@@
V-C Baselines
We have compared MM-MO with 5 state-of- the-art and classical merging methods, including Linear (Model Soups) [22], Task Arithmetic [23], TIES [27], DARE [17] and Model Breadcrumbs [53]. All detailed descriptions of the merging algorithms used for comparison are as follows:
Details of Parameter Settings for Baseline: We use the grid search method to determine the model merging configurations for the baselines, which is a commonly adopted practice in current model merging studies [17, 53]. Table II presents the parameter search ranges for each model merging method. We search the parameters of these baseline methods and select the optimal setting with the best performance. This approach aligns with previous work on model merging, as demonstrated in the DARE [17] paper. Specifically, Weight represents the weight proportion of each source model during merging, and Density denotes the sparsity level of each source model.
TABLE II: Grid searched ranges of parameters of each baseline.
Model Merging Methods Grid Search Ranges
Linear (Model Soups) Weight: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
Task Arithmetic Weight: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
TIES Weight: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
Density: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
DARE Weight: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
Density: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
Model Breadcrumbs Weight: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
Density: [0.1, 0.3, 0.5, 0.7, 0.9, 1.0]
V-D Evaluation Metrics
@@
VI Experimental Results and Analysis
The experimental sections of this paper are structured as follows.
In the main paper, we evaluate the general performance of MM-MO and provide detailed results on the BBH dataset (Subsection VI-A and VI-B). Next, we compare MM-MO with the SOTA evolutionary algorithm-based model merging method (Subsection VI-C). We also investigate the effects of different acquisition functions and varying model sizes and backbones for source models on MM-MO (Subsection VI-D and VI-E). Parameter sensitivity experiments are conducted to assess the impact of the number of evaluations (Subsection VI-F), followed by an ablation study of all modules within MM-MO (Subsection VI-G).
In the Appendix, we provide further supplementary experiments, including detailed results on the C-EVAL dataset (Appendix B-A), case studies on Mathematical Reasoning and Coding (Appendix B-B and B-C), an evaluation of MM-MO with a larger number of source models (Appendix B-D), and a comparative analysis of the validation sets used (Appendix B-E), offering a comprehensive evaluation of MM-MO’s performance and flexibility.
VI-A General Performance
TABLE III: Performance Comparison of different merging methods and single models. (Model size: 7B params)
@@ VI-A General Performance
DARE + TIES w/ MM-MO (Ours) Single Model 1 + 2 + 3 60.97 71.9 57.77 74.44 55.49 42.20 60.81 64.17
@@
VI-B Comparison on BBH Dataset
TABLE IV: Average accuracy over 27 BBH tasks.
@@ VI-B Comparison on BBH Dataset
DARE + TIES w/ MM-MO (Ours) Single Model 1 + 2 + 3 47.72
@@
TABLE V: Categorization of BBH tasks.
Category Tasks
Logical and Mathematical Reasoning logical_deduction_three_objects
logical_deduction_five_objects
logical_deduction_seven_objects
boolean_expressions
formal_fallacies
web_of_lies
causal_judgement
dyck_languages
multistep_arithmetic_two
object_counting
Language Understanding and Processing disambiguation_qa
hyperbaton
snarks
ruin_names
salient_translation_error_detection
word_sorting
Spatial and Visual Reasoning tracking_shuffled_objects_three_objects
tracking_shuffled_objects_five_objects
tracking_shuffled_objects_seven_objects
navigate
geometric_shapes
reasoning_about_colored_objects
Temporal and Date Understanding temporal_sequences
date_understanding
Common Sense and World Knowledge Application movie_recommendation
sports_understanding
penguins_in_a_table
@@
To validate our methods on difficult tasks, we utilize BBH, which includes a suite of 27 challenging BIG-Bench tasks that require multi-step reasoning. Table IV shows the performance of single models and the MM-MO merged model on the BBH datasets. Our MM-MO merged model achieves an average BBH score of 47.72, significantly outperforming each of the single models. This result highlights the effectiveness of our model merging strategy in enhancing overall performance.
@@
To further illustrate the advantages of our MM-MO method, we provide a detailed analysis of the performance across various sub-tasks within the BBH dataset. Fig. 4 visualizes the scores of single models and the MM-MO merged model on different types of sub-tasks, As can be seen from the radar chart, MM-MO (red) covers the largest area and has better performance. Notably, in the Logic and Mathematical Reasoning tasks, MM-MO achieves a score of 49.43, outperforming the best individual model, Single Model 2, which scored 45.87. In the Language Comprehension and Processing tasks, MM-MO reaches 46.77, significantly exceeding Single Model 3’s score of 42.72. Furthermore, in the Time and Date Understanding tasks, MM-MO scores 55.80, surpassing Single Model 1’s 51.00. This demonstrates that MM-MO effectively integrates the strengths of all individual models, resulting in a merged model that achieves remarkable performance across various tasks. In Table V, we provide a detailed breakdown of the tasks associated with each category.
VI-C MM-MO v.s. Evolutionary-model-merge (EMM)
TABLE VI: Performance Comparison of MM-MO and EMM. (Model size: 7B & 13B params)
Merging Method Models C-EVAL MMLU GSM8K Human Eval MBPP
Single Model 1 Qwen1.5-7B-Chat 68.7 60.06 54.59 46.95 34.20
Single Model 2 Liberated-Qwen1.5-7B 69.7 58.84 53.30 48.78 38.80
Single Model 3 firefly-qwen1.5-en-7b 70.0 51.66 49.81 33.54 28.40
Single Model 4 WizardLM-13B 34.8 51.47 55.50 35.98 30.60
Single Model 5 WizardMath-13B 30.2 51.27 60.50 14.02 25.20
Single Model 6 llama-2-13b-code-alpaca 33.2 52.99 29.72 21.95 30.00
DARE + TIES Single Model 1 + 2 + 3 69.5 60.06 55.72 49.39 37.80
DARE + TIES w/ EMM Single Model 1 + 2 + 3 68.0 58.99 62.02 34.76 29.60
DARE + TIES w/ MM-MO (Ours) Single Model 1 + 2 + 3 71.9 60.81 57.77 55.49 42.20
DARE + TIES Single Model 4 + 5 + 6 37.7 55.57 60.73 33.54 33.00
DARE + TIES w/ EMM Single Model 4 + 5 + 6 32.5 52.16 60.05 34.15 25.20
DARE + TIES w/ MM-MO (Ours) Single Model 4 + 5 + 6 38.0 56.36 62.85 36.59 36.80
In this section, we conduct a comparative experiment between our proposed method, MM-MO, and the Evolutionary-model-merge (EMM) approach introduced by Akiba et al. [11]. EMM models the configuration search for model merging as a single-objective optimization problem, utilizing the Covariance matrix adaptation evolution strategy (CMA-ES) [54] to identify the better merging configuration. Since CMA-ES is a single-objective evolutionary algorithm, we set the objective as the average score of the merged model across all tasks to enable a fair comparison. Both MM-MO and EMM were evaluated 30 real evaluation rounds.
As shown in Table VI, we compare the performance of MM-MO and EMM across 5 different tasks (following Qwen [3], the test questions cover four domains: Chinese, English, Mathematics, and Coding). The merged model obtained through our proposed MM-MO method achieves superior performance across all tasks compared to original DARE + TIES. Additionally, the MM-MO-optimized merged model outperforms all individual expert models on these five diverse tasks, demonstrating that modeling the model merging configuration search as a multi-objective optimization problem effectively yields better merging configurations. This contrasts with the single-objective approach in EMM, which fails to effectively balance multiple objectives. While the EMM-optimized merged model attains a higher score on the GSM8K task, its performance on the remaining tasks falls short of the original DARE + TIES.
Additionally, as model size increases, the advantages of MM-MO over EMM become more pronounced. With the model collection featuring larger model sizes (WizardLM-13B, WizardMath-13B, and Llama-2-13B-Code-Alpaca), MM-MO consistently outperforms EMM across all tasks. This result suggests that MM-MO’s enhanced acquisition strategy enables it to identify better configurations under limited number of real evaluations, compared to EMM. Furthermore, experimental results indicate that MM-MO demonstrates strong adaptability to larger model sizes and various model architectures.
Figure 5: Time overhead comparison between MM-MO and EMM under non-parallel evaluation.
Fig. 5 presents the time overhead comparison between EMM and MM-MO under non-parallel evaluation. Considering the performance improvements achieved by MM-MO over EMM, the additional experimental overhead remains within an acceptable range. Furthermore, as indicated by the results in Table VI, MM-MO can identify configurations with superior overall performance using fewer real evaluations, making it particularly advantageous in practical applications. Additionally, because evaluations of multiple new configurations can be conducted in parallel, the runtime for both EMM and MM-MO is amortizable in real-world scenarios.
VI-D MM-MO with Different Acquisition Function
TABLE VII: Performance Comparison of Different Acquisition Functions
Method Average Score C-EVAL MMLU GSM8K Human Eval MBPP
DARE + TIES 54.49 69.5 60.06 55.72 49.39 37.80
MM-MO / qNEHVI 56.09 71.6 61.07 57.16 50.61 40.00
MM-MO / qNParEGO 55.09 70.3 60.45 59.74 46.95 38.00
MM-MO / qEHVI (Ours) 57.63 71.9 60.81 57.77 55.49 42.20
To assess the impact of different acquisition functions on the performance of MM-MO, we compare qEHVI [30] with two other widely used acquisition functions: qNEHVI [55] and qParEGO [30].
The results in Table VII demonstrate that qEHVI outperforms the other acquisition functions. Acquisition function qEHVI ensures that each new candidate configuration contributes meaningfully to the Pareto front’s overall performance, thereby yielding balanced and comprehensive improvements across multiple objectives. In contrast, qNEHVI, although also designed to maximize hypervolume improvement, introduces additional computational complexity by modeling noise in the objective functions. This noise-handling feature, while beneficial in highly uncertain environments, can limit optimization efficiency in scenarios where objective evaluations are relatively stable, as is often the case with MM-MO. Meanwhile, qParEGO transforms the multi-objective optimization into a single-objective problem using scalarization, which simplifies computation but risks oversimplifying complex objective interactions. This linear decomposition may limit qParEGO’s ability to capture the trade-offs inherent in MM-MO’s multi-objective landscape. Overall, qEHVI’s direct optimization of the Pareto front’s hypervolume makes it particularly effective for multi-objective optimization in MM-MO, as it inherently balances improvements across objectives without the need for noise handling or decomposition, thereby leading to better performance.
VI-E MM-MO with Different Model Size and Source Models
@@
TABLE VIII: Performance comparison of different merging methods and single models. (Model size: 1.5B params)
@@
Single Model 7 Qwen2-1.5B-Instruct 47.05 68.2 62.09 40.77 42.68 30.00 54.58 42.70 35.37
Single Model 8 SauerkrautLM-1.5b 35.08 38.9 43.44 46.54 23.78 0.00 50.52 51.30 26.18
Single Model 9 Qwen2-1.5B-Ita 46.10 66.9 49.51 48.48 39.63 32.40 53.31 52.41 26.17
DARE + TIES w/ MM-MO (Ours) Single Model 7 + 8 + 9 51.04 69.3 63.15 50.65 43.90 35.60 54.84 54.54 36.33
To validate the versatility and generalization capabilities of our method, we applied MM-MO for model merging on source models with different model sizes and different Pre-Trained Backbones, followed by extensive evaluations of the merged models. As shown in Table VIII, the models obtained using MM-MO consistently outperformed any single model across all task types.
Furthermore, we selected source models with different model sizes and pre-training backbones, which is distinct from our previous experimental setup. This selection includes models optimized for specific domains and tasks. MM-MO demonstrated strong adaptability in merging models with these diverse characteristics. This adaptability highlights MM-MO’s robustness across models of varying complexity and suggesting significant potential for cross-domain model merging in diverse applications.
VI-F Parameter Sensitivity Experiment
TABLE IX: Performance of MM-MO with different number of function evaluations.
Single Model 1 Qwen1.5-7B-Chat 55.34 68.7 51.1 54.59 46.95
Single Model 2 Liberated-Qwen1.5-7B 55.47 69.7 50.1 53.30 48.78
Single Model 3 firefly-qwen1.5-en-7b 51.01 70.0 50.7 49.81 33.54
MM-MO / FE = 20 Single Model 1 + 2 + 3 54.62 70.7 51.5 53.60 42.68
MM-MO / FE = 40 Single Model 1 + 2 + 3 59.27 72.0 52.8 57.62 54.66
MM-MO / FE = 30 (Default) Single Model 1 + 2 + 3 59.47 71.9 52.7 57.77 55.49
In this section, we compare the performance of MM-MO with different number of function evaluations to validate the sensitivity of our proposed method to this parameter. Table IX illustrates the impact of varying number of function evaluations on the performance of MM-MO.
VI-G Ablation Study
TABLE X: Ablation study results.